The EXT2FS Library

The EXT2FS Library Version 1.44.5 December 2018

by Theodore Ts’o

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1997, 1998, 1999, 2000, 2001, 2002, 2003, 2004, 2005, 2006, 2007, 2008, 2009, 2010, 2011, 2012, 2013, 2014 Theodore Ts’o

Permission is granted to make and distribute verbatim copies of this manual provided the copyright notice and this permission notice are preserved on all copies.

Permission is granted to copy and distribute modified versions of this manual under the conditions for verbatim copying, provided that the entire resulting derived work is distributed under the terms of a permission notice identical to this one.

Permission is granted to copy and distribute translations of this manual into another language, under the above conditions for modified versions, except that this permission notice may be stated in a translation approved by the Foundation.

The EXT2FS Library 1

The EXT2FS Library

This manual documents the EXT2FS Library, version 1.44.5.

2 The EXT2FS Library (version 1.44.5)

1 Introduction to the EXT2FS Library

The EXT2FS library is designed to allow user-level programs to manipulate an ext2 filesystem.

2 EXT2FS Library Functions

2.1 Filesystem-level functions

The following functions operate on a filesystem handle. Most EXT2FS Library functions require a filesystem handle as their first argument. There are two functions which create a

filesystem handle, ext2fs_open and ext2fs_initialize.

The filesystem can also be closed using ext2fs_close, and any changes to the superblock and group descriptors can be written out to disk using ext2fs_flush.

2.1.1 Opening an ext2 filesystem

Most libext2fs functions take a filesystem handle of type ext2_filsys. A filesystem handle is created either by opening an existing filesystem using ext2fs_open, or by initializing a new filesystem using ext2fs_initialize.

[Function] errcode_t ext2fs_open (const char *name, int flags, int
   superblock, int block_size, io manager manager, ext2 filsys *ret_fs)

Opens a filesystem named name, using the the io manager manager to define the input/output routines needed to read and write the filesystem. In the case of the unix_io io manager, name is interpreted as the Unix filename of the filesystem image. This is often a device file, such as ‘/dev/hda1’.

The superblock parameter specifies the block number of the superblock which should be used when opening the filesystem. If superblock is zero, ext2fs_open will use the primary superblock located at offset 1024 bytes from the start of the filesystem image.

The block size parameter specifies the block size used by the filesystem. Normally this is determined automatically from the filesystem superblock. If block size is nonzero, it must match the block size found in the superblock, or the error EXT2_ET_ UNEXPECTED_BLOCK_SIZE will be returned. The block size parameter is also used to help find the superblock when superblock is non-zero.

The flags argument contains a bitmask of flags which control how the filesystem open should be handled.

EXT2_FLAG_RW

Open the filesystem for reading and writing. Without this flag, the filesystem is opened for reading only.

EXT2_FLAG_FORCE

Open the filesystem regardless of the feature sets listed in the superblock.

2.1.2 Closing and flushing out changes

[Function] errcode_t ext2fs_flush (ext2 filsys fs)

Write any changes to the high-level filesystem data structures in the fs filesystem. The following data structures will be written out:

• The filesystem superblock

• The filesystem group descriptors

• The filesystem bitmaps, if read in via ext2fs_read_bitmaps.

[Function] void ext2fs_free (ext2 filsys fs)

Close the io manager abstraction for fs and release all memory associated with the filesystem handle.

[Function] errcode_t ext2fs_close (ext2 filsys fs)

Flush out any changes to the high-level filesystem data structures using ext2fs_flush if the filesystem is marked dirty; then close and free the filesystem using ext2fs_free.

2.1.3 Initializing a filesystem

An ext2 filesystem is initializing by the mke2fs program. The two functions described here,

ext2fs_initialize and ext2fs_allocate_tables do much of the initial work for setting

up a filesystem. However, they don’t do the whole job. mke2fs calls ext2fs_initialize to set up the filesystem superblock, and calls ext2fs_allocate_tables to allocate space for the inode table, and the inode and block bitmaps. In addition, mke2fs must also initialize the inode tables by clearing them with zeros, create the root and lost+found directories, and reserve the reserved inodes.

[Function] errcode_t ext2fs_initialize (const char *name, int flags, struct

ext2 super block *param, io manager manager, ext2 filsys *ret_fs) This function is used by the mke2fs program to initialize a filesystem. The ext2fs_ initialize function creates a filesystem handle which is returned in ret fs that has been properly setup for a filesystem to be located in name, using the io manager manager. The prototype superblock in param is used to supply parameters such as the number of blocks in the filesystem, the block size, etc.

The ext2fs_initialize function does not actually do any I/O; that will be done when the application program calls ext2fs_close or ext2fs_flush. Also, this function only initializes the superblock and group descriptor structures. It does not create the inode table or the root directory. This must be done by the calling application, such as mke2fs.

The following values may be set in the param prototype superblock; if a value of 0 is found in a field, ext2fs_initialize will use a default value. The calling application should zero out the prototype entire superblock, and then fill in any appropriate values.

s_blocks_count

The number of blocks in the filesystem. This parameter is mandatory and must be set by the calling application.

s_inodes_count

The number of inodes in the filesystem. The default value is determined by calculating the size of the filesystem, and creating one inode for every 4096 bytes.

s_r_blocks_count

The number of blocks which should be reserved for the superuser. The default value is zero blocks.

s_log_block_size

The blocksize of the filesystem. Valid values are 0 (1024 bytes), 1 (2048 bytes), or 2 (4096 bytes). The default blocksize is 1024 bytes.

s_log_frag_size

The size of fragments. The ext2 filesystem does not support fragments (and may never support fragments). Currently this field must be the

same as s_log_block_size.
s_first_data_block

The first data block for the filesystem. For filesystem with a blocksize of 1024 bytes, this value must be at least 1, since the superblock is located in block number 1. For filesystems with larger blocksizes, the superblock is still located at an offset of 1024 bytes, so the superblock is located in block number 0. By default, this value is set to 1 for filesystems with a block size of 1024 bytes, or 0 for filesystems with larger blocksizes.

s_max_mnt_count

This field defines the number of times that the filesystem can be mounted before it should be checked using e2fsck. When e2fsck is run without the ‘-f’ option, e2fsck will skip the filesystem check if the number of times that the filesystem has been mounted is less than s_max_mnt_count and if the interval between the last time a filesystem check was performed and the current time is less than s_checkinterval (see below). The default value of s_max_mnt_count is 20.

s_checkinterval

This field defines the minimal interval between filesystem checks. See the previous entry for a discussion of how this field is used by e2fsck. The default value of this field is 180 days (six months).

s_errors This field defines the behavior which should be used by the kernel of errors are detected in the filesystem. Possible values include:

‘EXT2_ERRORS_CONTINUE’

Continue execution when errors are detected.

‘EXT2_ERRORS_RO’

Remount the filesystem read-only.

‘EXT2_ERRORS_PANIC’

Panic.

The default behavior is ‘EXT2_ERRORS_CONTINUE’.

[Function] errcode_t ext2fs_allocate_tables (ext2 filsys fs)

Allocate space for the inode table and the block and inode bitmaps. The inode tables and block and inode bitmaps aren’t actually initialized; this function just allocates the space for them.

2.1.4 Filesystem flag functions

The filesystem handle has a number of flags which can be manipulated using the following function. Some of these flags affect how the libext2fs filesystem behaves; others are provided solely for the application’s convenience.

[Function] void ext2fs_mark_changed (ext2 filsys fs)
[Function] int ext2fs_test_changed (ext2 filsys fs)

This flag indicates whether or not the filesystem has been changed. It is not used by the ext2fs library.

[Function] void ext2fs_mark_super_dirty (ext2 filsys fs)

Mark the filesystem fs as being dirty; this will cause the superblock information to be

flushed out when ext2fs_close is called. ext2fs_mark_super_dirty will also set

the filesystem changed flag. The dirty flag is automatically cleared by ext2fs_flush when the superblock is written to disk.

[Function] void ext2fs_mark_valid (ext2 filsys fs)
[Function] void ext2fs_unmark_valid (ext2 filsys fs)
[Function] int ext2fs_test_valid (ext2 filsys fs)

This flag indicates whether or not the filesystem is free of errors. It is not used by libext2fs, and is solely for the application’s convenience.

[Function] void ext2fs_mark_ib_dirty (ext2 filsys fs)
[Function] void ext2fs_mark_bb_dirty (ext2 filsys fs)
[Function] int ext2fs_test_ib_dirty (ext2 filsys fs)
[Function] int ext2fs_test_bb_dirty (ext2 filsys fs)

These flags indicate whether or not the inode or block bitmaps have been modified. If the flag is set, it will cause the appropriate bitmap to be written when the filesystem is closed or flushed.

2.2 File I/O Functions

The following functions provide a convenient abstraction to read or write a file in an filesystem. The interface is similar in spirit to the Linux/POSIX file I/O system calls.

2.2.1 File handle manipulation

The file handle functions much like a file descriptor in the Linux/POSIX file I/O system calls. Unlike the Linux/POSIX system calls, files are opened via inode numbers instead of via pathnames. To resolve a pathname to an inode number, use the function ext2fs_namei

or to create a new file, use ext2fs_new_inode and ext2fs_link.

[Function] errcode_t ext2fs_file_open2 (ext2 filsys fs, ext2 ino t ino, struct

ext2 inode *inode, int flags, ext2 file t *ret)

[Function] errcode_t ext2fs_file_open (ext2 filsys fs, ext2 ino t ino, int

flags, ext2 file t *ret) Opens a file identified by inode number ino in filesystem fs and returns a file handle in ret. If an inode structure is provided in inode, then it is used instead of reading the inode from the filesystem.

The flags argument contains a bitmask of flags which control how the file should be opened.

EXT2_FILE_WRITE

Open the file for reading and writing. Without this flag, the file is opened for writing only.

EXT2_FILE_CREATE

Create the file if it is not already present.

[Function] ext2_filsys ext2fs_file_get_fs (ext2 file t file)

Return the filesystem handle where the open file file was opened.

[Function] errcode_t ext2fs_file_close (ext2 file t file)

Close the file handle file.

[Function] errcode_t ext2fs_file_flush (ext2 file t file)

Force any data written via ext2fs_file_write to disk.

2.2.2 Reading and writing data

[Function] errcode_t ext2fs_file_read (ext2 file t file, void *buf, unsigned

int wanted, unsigned int *got) Read wanted bytes of data from file store it in the buffer buf. The number of bytes that was actually read is returned via got.

[Function] errcode_t ext2fs_file_write (ext2 file t file, const void *buf,

unsigned int nbytes, unsigned int *written) Write wanted bytes of data from the buffer buf to the current file position of file. The number of bytes that was actually written is returned via written.

2.2.3 Changing the file offset

[Function] errcode_t ext2fs_file_llseek (ext2 file t file, u64 offset, int
   whence, u64 *ret_pos)
[Function] errcode_t ext2fs_file_lseek (ext2 file t file, ext2 off t offset,
   int whence, ext2 off t *ret_pos)

Change the current file position of file according to the directive whence as follows:

EXT2_SEEK_SET

The file position is set to offset bytes from the beginning of the file.

EXT2_SEEK_CUR

The file position set to its current location plus offset bytes.

EXT2_SEEK_END

The file position is set to the size of the file plus offset bytes.

The current offset is returned via ret pos.

2.2.4 Getting the file size

[Function] errcode_t ext2fs_file_get_lsize (ext2 file t file, u64
   *ret_size)

Return the size of the file file in ret size.

[Function] ext2_off_t ext2fs_file_get_size (ext2 file t file)

Return the size of the file file.

2.3 Inode Functions

2.3.1 Reading and writing inodes

[Function] errcode_t ext2fs_read_inode (ext2 filsys fs, ext2 ino t ino, struct

ext2 inode *inode) Read the inode number ino into inode.

[Function] errcode_t ext2fs_write_inode (ext2 filsys fs, ext2 ino t ino, struct

ext2 inode *inode) Write inode to inode ino.

2.3.2 Iterating over inodes in a filesystem

The inode scan abstraction is useful for iterating over all the inodes in a filesystem.

[Function] errcode_t ext2fs_open_inode_scan (ext2 filsys fs, int
   buffer_blocks, ext2 inode scan *scan)

Initialize the iteration variable scan. This variable is used by ext2fs_get_next_ inode. The buffer blocks parameter controls how many blocks of the inode table are read in at a time. A large number of blocks requires more memory, but reduces the overhead in seeking and reading from the disk. If buffer blocks is zero, a suitable default value will be used.

[Function] void ext2fs_close_inode_scan (ext2 inode scan scan)

Release the memory associated with scan and invalidate it.

[Function] errcode_t ext2fs_get_next_inode (ext2 inode scan scan, ext2 ino t

*ino, struct ext2 inode *inode) This function returns the next inode from the filesystem; the inode number of the inode is stored in ino, and the inode is stored in inode.

If the inode is located in a block that has been marked as bad, ext2fs_get_next_

inode will return the error EXT2_ET_BAD_BLOCK_IN_INODE_TABLE.
[Function] errcode_t ext2fs_inode_scan_goto_blockgroup (ext2 inode scan
   scan, int group)

Start the inode scan at a particular ext2 blockgroup, group. This function may be safely called at any time while scan is valid.

[Function] void ext2fs_set_inode_callback (ext2 inode scan scan, errcode t

(*done group)(ext2 filsys fs, ext2 inode scan scan, dgrp t group, void *

   private), void *done_group_data)

Register a callback function which will be called by ext2_get_next_inode when all of the inodes in a block group have been processed.

[Function] int ext2fs_inode_scan_flags (ext2 inode scan scan, int
   set_flags, int clear_flags)

Set the scan flags set flags and clear the scan flags clear flags. The following flags can be set using this interface:

‘EXT2_SF_SKIP_MISSING_ITABLE’

When a block group is missing an inode table, skip it. If this flag is not set ext2fs_get_next_inode will return the error EXT2 ET MISSING INODE TABLE.

2.3.3 Iterating over blocks in an inode

[Function] errcode_t ext2fs_block_iterate (ext2 filsys fs, ext2 ino t ino, int

flags, char *block buf, int (*func)(ext2 filsys fs, blk t *blocknr, int

   blockcnt, void *private), void *private)

Iterate over all of the blocks in inode number ino in filesystem fs, by calling the function func for each block in the inode. The block buf parameter should either be NULL, or if the ext2fs_block_iterate function is called repeatedly, the overhead of allocating and freeing scratch memory can be avoided by passing a pointer to a scratch buffer which must be at least as big as three times the filesystem’s blocksize.

The flags parameter controls how the iterator will function:

‘BLOCK_FLAG_HOLE’

This flag indicates that the interator function should be called on blocks where the block number is zero (also known as “holes”.) It is also known as BLOCK FLAG APPEND, since it is also used by functions such as ext2fs expand dir() to add a new block to an inode.

‘BLOCK_FLAG_DEPTH_TRAVERSE’

This flag indicates that the iterator function for the indirect, doubly indirect, etc. blocks should be called after all of the blocks contained in the indirect blocks are processed. This is useful if you are going to be deallocating blocks from an inode.

‘BLOCK_FLAG_DATA_ONLY’

This flag indicates that the iterator function should be called for data blocks only.

The callback function func is called with a number of parameters; the fs and private parameters are self-explanatory, and their values are taken from the parameters to ext2fs_block_iterate. (The private data structure is generally used by callers to ext2fs_block_iterate so that some private data structure can be passed to the callback function. The blockcnt parameter, if non-negative, indicates the logical block number of a data block in the inode. If blockcnt is less than zero, then func was called on a metadata block, and blockcnt will be one of the following values: BLOCK COUNT IND, BLOCK COUNT DIND, BLOCK COUNT TIND, or BLOCK COUNT TRANSLATOR. The blocknr is a pointer to the inode or indirect block entry listing physical block number. The callback function may modify the physical block number, if it returns the BLOCK CHANGED flag.

The callback function func returns a result code which is composed of the logical OR of the following flags:

‘BLOCK_CHANGED’

This flag indicates that callback function has modified the physical block number pointed to by blocknr.

‘BLOCK_ABORT’

This flag requests that ext2fs_block_iterate to stop immediately and return to the caller.

[Function] errcode_t ext2fs_block_iterate2 (ext2 filsys fs, ext2 ino t ino, int

flags, char *block buf, int (*func)(ext2 filsys fs, blk t *blocknr,

   e2 blkcnt t blockcnt, blk t ref_blk, int ref_offset, void *private), void
   *private)

This function is much like ext2fs_block_iterate, except that the blockcnt type is a 64-bit signed quantity, to support larger files, and the addition of the ref blk and ref offset arguments passed to the callback function, which identify the location of the physical block pointed to by pointer blocknr. If ref blk is zero, then ref offset contains the offset into the i_blocks array. If ref blk is non-zero, then the physical block location is contained inside an indirect block group, and ref offset contains the offset into the indirect block.

2.3.4 Convenience functions for Inodes

[Function] errcode_t ext2fs_get_blocks (ext2 filsys fs, ext2 ino t ino, blk t

   *blocks)

Returns an array of blocks corresponding to the direct, indirect, doubly indirect, and triply indirect blocks as stored in the inode structure.

[Function] errcode_t ext2fs_check_directory (ext2 filsys fs, ext2 ino t ino)

Returns 0 if ino is a directory, and ENOTDIR if it is not.

[Function] int ext2fs_inode_has_valid_blocks (struct ext2 inode *inode)

Returns 1 if the inode’s block entries actually valid block entries, and 0 if not. Inodes which represent devices and fast symbolic links do not contain valid block entries.

2.4 Directory functions

2.4.1 Directory block functions

[Function] errcode_t ext2fs_read_dir_block (ext2 filsys fs, blk t block, void
   *buf)

This function reads a directory block, performing byte swapping if necessary.

[Function] errcode_t ext2fs_write_dir_block (ext2 filsys fs, blk t block,

void *buf) This function writes a directory block, performing byte swapping if necessary.

[Function] errcode_t ext2fs_new_dir_block (ext2 filsys fs, ext2 ino t
   dir_ino, ext2 ino t parent_ino, char **block)

This function creates a new directory block in block. If dir ino is non-zero, then dir ino and parent ino are used to initialize directory entries for ‘.’ and ‘..’, respectively.

2.4.2 Iterating over a directory

[Function] errcode_t ext2fs_dir_iterate (ext2 filsys fs, ext2 ino t dir, int

flags, char *block_buf, int (*func)(struct ext2 dir entry *dirent, int

   offset, int blocksize, char *buf, void *private), void *private)

This function iterates over all of the directory entries in the directory dir, calling the callback function func for each directory entry in the directory. The block buf parameter should either be NULL, or if the ext2fs_dir_iterate function is called repeatedly, the overhead of allocating and freeing scratch memory can be avoided by passing a pointer to a scratch buffer which must be at least as big as the filesystem’s blocksize.

The flags parameter controls how the iterator will function:

‘DIRENT_FLAG_INCLUDE_EMPTY’

This flag indicates that the callback function should be called even for deleted or empty directory entries.

2.4.3 Creating and expanding directories

[Function] errcode_t ext2fs_mkdir (ext2 filsys fs, ext2 ino t parent, ext2 ino t

inum, const char *name) This function creates a new directory. If inum is zero, then a new inode will be allocated; otherwise, the directory will be created in the inode specified by inum. If name specifies the name of the new directory; if it is non-NULL, then the new directory will be linked into the parent directory parent.

[Function] errcode_t ext2fs_expand_dir (ext2 filsys fs, ext2 ino t dir)

This function adds a new empty directory block and appends it to the directory dir. This allows functions such as ext2fs_link to add new directory entries to a directory which is full.

2.4.4 Creating and removing directory entries

[Function] errcode_t ext2fs_link (ext2 filsys fs, ext2 ino t dir, const char

*name, ext2 ino t ino, int flags) This function adds a new directory entry to the directory dir, with name and ino specifying the name and inode number in the directory entry, respectively.

The low 3 bits of the flags field is used to specify the file type of inode: (No other flags are currently defined.)

‘EXT2_FT_UNKNOWN’

The file type is unknown.

‘EXT2_FT_REG_FILE’

The file type is a normal file.

‘EXT2_FT_DIR’

The file type is a directory.

‘EXT2_FT_CHRDEV’

The file type is a character device.

‘EXT2_FT_BLKDEV’

The file type is a block device.

‘EXT2_FT_FIFO’

The file type is a named pipe.

‘EXT2_FT_SOCK’

The file type is a unix domain socket.

‘EXT2_FT_SYMLINK’

The file type is a symbolic link.

[Function] errcode_t ext2fs_unlink (ext2 filsys fs, ext2 ino t dir, const char

*name, ext2 ino t ino, int flags) This function removes a directory entry from dir. The directory entry to be removed is the first one which is matched by name and ino. If name is non-NULL, the directory entry’s name must match name. If ino is non-zero, the directory entry’s inode number must match ino. No flags are currently defined for ext2fs_unlink; callers should pass in zero to this parameter.

2.4.5 Looking up filenames

[Function] errcode_t ext2fs_lookup (ext2 filsys fs, ext2 ino t dir, const char *name, int namelen, char *buf, ext2 ino t *inode)

[Function] errcode_t ext2fs_namei (ext2 filsys fs, ext2 ino t root, ext2 ino t cwd, const char *name, ext2 ino t *inode)

[Function] errcode_t ext2fs_namei_follow (ext2 filsys fs, ext2 ino t root,

ext2 ino t cwd, const char *name, ext2 ino t *inode)

[Function] errcode_t ext2fs_follow_link (ext2 filsys fs, ext2 ino t root,

ext2 ino t cwd, ext2 ino t inode, ext2 ino t *res inode)

2.4.6 Translating inode numbers to filenames

[Function] errcode_t ext2fs_get_pathname (ext2 filsys fs, ext2 ino t dir,

ext2 ino t ino, char **name)

2.5 Bitmap Functions

2.5.1 Reading and Writing Bitmaps

[Function] errcode_t ext2fs_write_inode_bitmap (ext2 filsys fs)
[Function] errcode_t ext2fs_write_block_bitmap (ext2 filsys fs)
[Function] errcode_t ext2fs_read_inode_bitmap (ext2 filsys fs)
[Function] errcode_t ext2fs_read_block_bitmap (ext2 filsys fs)
[Function] errcode_t ext2fs_read_bitmaps (ext2 filsys fs)
[Function] errcode_t ext2fs_write_bitmaps (ext2 filsys fs)

2.5.2 Allocating Bitmaps

[Function] errcode_t ext2fs_allocate_generic_bitmap ( u32 start, u32

end, u32 real_end, const char *descr, ext2fs generic bitmap *ret)

[Function] errcode_t ext2fs_allocate_block_bitmap (ext2 filsys fs, const

char *descr, ext2fs block bitmap *ret)

[Function] errcode_t ext2fs_allocate_inode_bitmap (ext2 filsys fs, const

char *descr, ext2fs inode bitmap *ret)

2.5.3 Freeing bitmaps

[Function] void ext2fs_free_generic_bitmap (ext2fs inode bitmap bitmap)
[Function] void ext2fs_free_block_bitmap (ext2fs block bitmap bitmap)
[Function] void ext2fs_free_inode_bitmap (ext2fs inode bitmap bitmap)

2.5.4 Bitmap Operations

[Function] void ext2fs_mark_block_bitmap (ext2fs block bitmap bitmap, blk t
        block)
[Function] void ext2fs_unmark_block_bitmap (ext2fs block bitmap bitmap,
        blk t block)
[Function] int ext2fs_test_block_bitmap (ext2fs block bitmap bitmap, blk t
        block)

These functions set, clear, and test bits in a block bitmap bitmap.

[Function] void ext2fs_mark_inode_bitmap (ext2fs inode bitmap bitmap,

ext2 ino t inode)

[Function] void ext2fs_unmark_inode_bitmap (ext2fs inode bitmap bitmap,

ext2 ino t inode)

[Function] int ext2fs_test_inode_bitmap (ext2fs inode bitmap bitmap,

ext2 ino t inode) These functions set, clear, and test bits in an inode bitmap bitmap.

[Function] void ext2fs_fast_mark_block_bitmap (ext2fs block bitmap bitmap,
        blk t block)
[Function] void ext2fs_fast_unmark_block_bitmap (ext2fs block bitmap
        bitmap, blk t block)
[Function] int ext2fs_fast_test_block_bitmap (ext2fs block bitmap bitmap,
        blk t block)
[Function] void ext2fs_fast_mark_inode_bitmap (ext2fs inode bitmap bitmap,

ext2 ino t inode)

[Function] void ext2fs_fast_unmark_inode_bitmap (ext2fs inode bitmap
        bitmap, ext2 ino t inode)
[Function] int ext2fs_fast_test_inode_bitmap (ext2fs inode bitmap bitmap,

ext2 ino t inode) These “fast” functions are like their normal counterparts; however, they are implemented as inline functions and do not perform bounds checks on the inode number or block number; they are assumed to be correct. They should only be used in speedcritical applications, where the inode or block number has already been validated by other means.

[Function] blk_t ext2fs_get_block_bitmap_start (ext2fs block bitmap
        bitmap)
[Function] ext2_ino_t ext2fs_get_inode_bitmap_start (ext2fs inode bitmap
        bitmap)

Return the first inode or block which is stored in the bitmap.

[Function] blk_t ext2fs_get_block_bitmap_end (ext2fs block bitmap bitmap)
[Function] ext2_ino_t ext2fs_get_inode_bitmap_end (ext2fs inode bitmap
        bitmap)

Return the last inode or block which is stored in the bitmap.

2.5.5 Comparing bitmaps

[Function] errcode_t ext2fs_compare_block_bitmap (ext2fs block bitmap bm1,

ext2fs block bitmap bm2)

[Function] errcode_t ext2fs_compare_inode_bitmap (ext2fs inode bitmap bm1,

ext2fs inode bitmap bm2)

2.5.6 Modifying Bitmaps

[Function] errcode_t ext2fs_fudge_inode_bitmap_end (ext2fs inode bitmap

bitmap, ext2 ino t end, ext2 ino t *oend)

[Function] errcode_t ext2fs_fudge_block_bitmap_end (ext2fs block bitmap
        bitmap, blk t end, blk t *oend)

2.5.7 Resizing Bitmaps

[Function] errcode_t ext2fs_resize_generic_bitmap ( u32 new_end, u32

new_real_end, ext2fs generic bitmap bmap)

[Function] errcode_t ext2fs_resize_inode_bitmap ( u32 new_end, u32

new_real_end, ext2fs inode bitmap bmap)

[Function] errcode_t ext2fs_resize_block_bitmap ( u32 new_end, u32

new_real_end, ext2fs block bitmap bmap)

2.5.8 Clearing Bitmaps

[Function] void ext2fs_clear_inode_bitmap (ext2fs inode bitmap bitmap)

This function sets all of the bits in the inode bitmap bitmap to be zero.

[Function] void ext2fs_clear_block_bitmap (ext2fs block bitmap bitmap)

This function sets all of the bits in the block bitmap bitmap to be zero.

2.6 EXT2 data abstractions

The ext2 library has a number of abstractions which are useful for ext2 utility programs.

2.6.1 Badblocks list management

[Function] errcode_t ext2fs_badblocks_list_create (ext2 badblocks list
        *ret, int size)
[Function] void ext2fs_badblocks_list_free (ext2 badblocks list bb)
[Function] errcode_t ext2fs_badblocks_list_add (ext2 badblocks list bb,

blk t blk)

[Function] int ext2fs_badblocks_list_test (ext2 badblocks list bb, blk t blk)
[Function] errcode_t ext2fs_badblocks_list_iterate_begin

(ext2 badblocks list bb, ext2 badblocks iterate *ret)

[Function] int ext2fs_badblocks_list_iterate (ext2 badblocks iterate iter,

blk t *blk)

[Function] void ext2fs_badblocks_list_iterate_end (ext2 badblocks iterate
        iter)
[Function] errcode_t ext2fs_update_bb_inode (ext2 filsys fs,

ext2 badblocks list bb_list)

[Function] errcode_t ext2fs_read_bb_inode (ext2 filsys fs, ext2 badblocks list

        *bb_list)
[Function] errcode_t ext2fs_read_bb_FILE (ext2 filsys fs, FILE *f,

ext2 badblocks list *bb_list, void (*invalid)(ext2 filsys fs, blk t blk))

2.6.2 Directory-block list management

The dblist abstraction stores a list of blocks belonging to directories. This list can be useful when a program needs to interate over all directory entries in a filesystem; e2fsck does this in pass 2 of its operations, and debugfs needs to do this when it is trying to turn an inode number into a pathname.

[Function] errcode_t ext2fs_init_dblist (ext2 filsys fs, ext2 dblist

        *ret_dblist)

Creates a dblist data structure and returns it in ret dblist.

[Function] void ext2fs_free_dblist (ext2 dblist dblist)

Free a dblist data structure.

[Function] errcode_t ext2fs_add_dir_block (ext2 dblist dblist, ext2 ino t
   ino, blk t blk, int blockcnt)

Add an entry to the dblist data structure. This call records the fact that block number blockcnt of directory inode ino is stored in block blk.

[Function] errcode_t ext2fs_set_dir_block (ext2 dblist dblist, ext2 ino t
   ino, blk t blk, int blockcnt)

Change an entry in the dblist data structure; this changes the location of block number blockcnt of directory inode ino to be block blk.

[Function] errcode_t ext2fs_dblist_iterate (ext2 dblist dblist, int

(*func)(ext2 filsys fs, struct ext2 db entry *db_info, void *private), void

   *private)

This iterator calls func for every entry in the dblist data structure.

[Function] errcode_t ext2fs_dblist_dir_iterate (ext2 dblist dblist, int

flags, char *block_buf, int (*func)(ext2 ino t dir, int entry, struct ext2 dir entry *dirent, int offset, int blocksize, char *buf, void

   *private), void *private)

This iterator takes reads in the directory block indicated in each dblist entry, and calls func for each directory entry in each directory block. If dblist contains all the directory blocks in a filesystem, this function provides a convenient way to iterate over all directory entries for that filesystem.

2.6.3 Inode count functions

The icount abstraction is a specialized data type used by e2fsck to store how many times a particular inode is referenced by the filesystem. This is used twice; once to store the actual number of times that the inode is reference; and once to store the claimed number of times the inode is referenced according to the inode structure.

This abstraction is designed to be extremely efficient for storing this sort of information, by taking advantage of the following properties of inode counts, namely (1) inode counts are very often zero (because the inode is currently not in use), and (2) many files have a inode count of 1 (because they are a file which has no additional hard links).

[Function] errcode_t ext2fs_create_icount2 (ext2 filsys fs, int flags, int

size, ext2 icount t hint, ext2 icount t *ret) Creates an icount structure for a filesystem fs, with initial space for size inodes whose count is greater than 1. The flags parameter is either 0 or EXT2_ICOUNT_ OPT_INCREMENT, which indicates that icount structure should be able to increment inode counts quickly. The icount structure is returned in ret. The returned icount structure initially has a count of zero for all inodes.

The hint parameter allows the caller to optionally pass in another icount structure which is used to initialize the array of inodes whose count is greater than 1. It is used purely as a speed optimization so that the icount structure can determine in advance which inodes are likely to contain a count grater than 1.

[Function] void ext2fs_free_icount (ext2 icount t icount)

Frees an icount structure.

[Function] errcode_t ext2fs_icount_fetch (ext2 icount t icount, ext2 ino t
   ino, u16 *ret)

Returns in ret the count for a particular inode ino.

[Function] errcode_t ext2fs_icount_increment (ext2 icount t icount,

ext2 ino t ino, u16 *ret) Increments the ref count for inode ino.

[Function] errcode_t ext2fs_icount_decrement (ext2 icount t icount,

ext2 ino t ino, u16 *ret) Decrements the ref count for inode ino.

[Function] errcode_t ext2fs_icount_store (ext2 icount t icount, ext2 ino t
   ino, u16 count)

Sets the reference count for inode ino to be count.

[Function] ext2_ino_t ext2fs_get_icount_size (ext2 icount t icount)

Returns the current number of inodes in icount which has a count greater than 1.

[Function] errcode_t ext2fs_icount_validate (ext2 icount t icount, FILE

*f) Validates the internal rep invariant of icount; if there are any problems, print out debugging information to f. This function is intended for debugging and testing use only.

2.7 Byte-swapping functions

[Function] void ext2fs_swap_super (struct ext2 super block * super)

[Function] void ext2fs_swap_group_desc (struct ext2 group desc *gdp)

[Function] void ext2fs_swap_inode (ext2 filsys fs, struct ext2 inode *to, struct ext2 inode *from, int hostorder)

[Function] int ext2fs_native_flag (void)

2.8 Other functions

/* alloc.c */

[Function] errcode_t ext2fs_new_inode (ext2 filsys fs, ext2 ino t dir, int mode,

ext2fs inode bitmap map, ext2 ino t *ret)

[Function] errcode_t ext2fs_new_block (ext2 filsys fs, blk t goal,

ext2fs block bitmap map, blk t *ret)

[Function] errcode_t ext2fs_get_free_blocks (ext2 filsys fs, blk t start,

blk t finish, int num, ext2fs block bitmap map, blk t *ret)

/* check desc.c */

[Function] errcode_t ext2fs_check_desc (ext2 filsys fs)
[Function] errcode_t ext2fs_get_num_dirs (ext2 filsys fs, ext2 ino t
   *ret_num_dirs)

/* getsize.c */

[Function] errcode_t ext2fs_get_device_size (const char *file, int
   blocksize, blk t *retblocks)

/* ismounted.c */

[Function] errcode_t ext2fs_check_if_mounted (const char *file, int
   *mount_flags)

/* version.c */

[Function] int ext2fs_get_library_version (const char **ver_string, const
   char **date_string)

This function returns the current version of the ext2 library. The return value contains an integer version code, which consists of the major version number of the library multiplied by 100, plus the minor version number of the library. Hence, if the library version is 1.08, the returned value will be 108.

If ver string and/or date string are non-NULL, they will be set to point at a constant string containing the library version and/or release date, respectively.

[Function] int ext2fs_parse_version_string (const char *ver_string)

This function takes a version string which may included in an application and returns a version code using the same algorithm used by ext2fs_get_library_version. It can be used by programs included in the e2fsprogs distribution to assure that they are using an up-to-date ext2 shared library.

/* inline functions */

[Function] int ext2fs_group_of_blk (ext2 filsys fs, blk t blk)

This function returns the block group which contains the block blk.

[Function] int ext2fs_group_of_ino (ext2 filsys fs, ext2 ino t ino)

This function returns the block group which contains the inode ino.

Concept Index 19

Concept Index

(Index is nonexistent)

20 The EXT2FS Library (version 1.44.5)

Function and Type Index

ext2fs_add_dir_block. . . . . . . . . . . . . . . . . . . . . . . 16

ext2fs_allocate_block_bitmap . . . . . . . . . . . . . . 13
ext2fs_allocate_generic_bitmap . . . . . . . . . . . . 13
ext2fs_allocate_inode_bitmap . . . . . . . . . . . . . . 13
ext2fs_allocate_tables . . . . . . . . . . . . . . . . . . . . . 5
ext2fs_badblocks_list_add . . . . . . . . . . . . . . . . . 15
ext2fs_badblocks_list_create . . . . . . . . . . . . . . 15
ext2fs_badblocks_list_free . . . . . . . . . . . . . . . . 15
ext2fs_badblocks_list_iterate . . . . . . . . . . . . . 15
ext2fs_badblocks_list_iterate_begin . . . . . . 15
ext2fs_badblocks_list_iterate_end. . . . . . . . . 15
ext2fs_badblocks_list_test . . . . . . . . . . . . . . . . 15

ext2fs_block_iterate. . . . . . . . . . . . . . . . . . . . . . . . 9 ext2fs_block_iterate2 . . . . . . . . . . . . . . . . . . . . . 10 ext2fs_check_desc . . . . . . . . . . . . . . . . . . . . . . . . . . 18

ext2fs_check_directory . . . . . . . . . . . . . . . . . . . . 10
ext2fs_check_if_mounted . . . . . . . . . . . . . . . . . . . 18
ext2fs_clear_block_bitmap . . . . . . . . . . . . . . . . . 15
ext2fs_clear_inode_bitmap . . . . . . . . . . . . . . . . . 15

ext2fs_close . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

ext2fs_close_inode_scan . . . . . . . . . . . . . . . . . . . . 8
ext2fs_compare_block_bitmap . . . . . . . . . . . . . . . 14
ext2fs_compare_inode_bitmap . . . . . . . . . . . . . . . 14

ext2fs_create_icount2 . . . . . . . . . . . . . . . . . . . . . 16

ext2fs_dblist_dir_iterate . . . . . . . . . . . . . . . . . 16

ext2fs_dblist_iterate . . . . . . . . . . . . . . . . . . . . . 16 ext2fs_dir_iterate . . . . . . . . . . . . . . . . . . . . . . . . . 11 ext2fs_expand_dir . . . . . . . . . . . . . . . . . . . . . . . . . . 11

ext2fs_fast_mark_block_bitmap . . . . . . . . . . . . . 14
ext2fs_fast_mark_inode_bitmap . . . . . . . . . . . . . 14
ext2fs_fast_test_block_bitmap . . . . . . . . . . . . . 14
ext2fs_fast_test_inode_bitmap . . . . . . . . . . . . . 14
ext2fs_fast_unmark_block_bitmap . . . . . . . . . . . 14
ext2fs_fast_unmark_inode_bitmap . . . . . . . . . . . 14

ext2fs_file_close . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 ext2fs_file_flush . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 ext2fs_file_get_fs . . . . . . . . . . . . . . . . . . . . . . . . . . 7 ext2fs_file_get_lsize. . . . . . . . . . . . . . . . . . . . . . . 8 ext2fs_file_get_size. . . . . . . . . . . . . . . . . . . . . . . . 8 ext2fs_file_llseek . . . . . . . . . . . . . . . . . . . . . . . . . . 7 ext2fs_file_lseek . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 ext2fs_file_open . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 ext2fs_file_open2 . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 ext2fs_file_read . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 ext2fs_file_write . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 ext2fs_flush . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 ext2fs_follow_link . . . . . . . . . . . . . . . . . . . . . . . . . 12 ext2fs_free . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

ext2fs_free_block_bitmap . . . . . . . . . . . . . . . . . . 13

ext2fs_free_dblist . . . . . . . . . . . . . . . . . . . . . . . . . 16

ext2fs_free_generic_bitmap . . . . . . . . . . . . . . . . 13

ext2fs_free_icount . . . . . . . . . . . . . . . . . . . . . . . . . 17

ext2fs_free_inode_bitmap . . . . . . . . . . . . . . . . . . 13
ext2fs_fudge_block_bitmap_end . . . . . . . . . . . . . 14

                                                ext2fs_fudge_inode_bitmap_end . . . . . . . . . . . . . 14
                                                ext2fs_get_block_bitmap_end . . . . . . . . . . . . . . . 14
                                                ext2fs_get_block_bitmap_start . . . . . . . . . . . . . 14

ext2fs_get_blocks . . . . . . . . . . . . . . . . . . . . . . . . . . 10

                                                ext2fs_get_device_size . . . . . . . . . . . . . . . . . . . . 18
                                                ext2fs_get_free_blocks . . . . . . . . . . . . . . . . . . . . 18
                                                ext2fs_get_icount_size . . . . . . . . . . . . . . . . . . . . 17
                                                ext2fs_get_inode_bitmap_end . . . . . . . . . . . . . . . 14
                                                ext2fs_get_inode_bitmap_start . . . . . . . . . . . . . 14
                                                ext2fs_get_library_version . . . . . . . . . . . . . . . . 18

ext2fs_get_next_inode. . . . . . . . . . . . . . . . . . . . . . . 8 ext2fs_get_num_dirs. . . . . . . . . . . . . . . . . . . . . . . . 18 ext2fs_get_pathname. . . . . . . . . . . . . . . . . . . . . . . . 13 ext2fs_group_of_blk. . . . . . . . . . . . . . . . . . . . . . . . 18 ext2fs_group_of_ino. . . . . . . . . . . . . . . . . . . . . . . . 18

                                                ext2fs_icount_decrement . . . . . . . . . . . . . . . . . . . 17

ext2fs_icount_fetch. . . . . . . . . . . . . . . . . . . . . . . . 17

                                                ext2fs_icount_increment . . . . . . . . . . . . . . . . . . . 17

ext2fs_icount_store. . . . . . . . . . . . . . . . . . . . . . . . 17

                                                ext2fs_icount_validate . . . . . . . . . . . . . . . . . . . . 17

ext2fs_init_dblist . . . . . . . . . . . . . . . . . . . . . . . . . 15 ext2fs_initialize . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

                                                ext2fs_inode_has_valid_blocks . . . . . . . . . . . . . 10
                                                ext2fs_inode_scan_flags . . . . . . . . . . . . . . . . . . . . 9
                                                ext2fs_inode_scan_goto_blockgroup. . . . . . . . . . 8

ext2fs_link . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 ext2fs_lookup . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 ext2fs_mark_bb_dirty. . . . . . . . . . . . . . . . . . . . . . . . 6

                                                ext2fs_mark_block_bitmap . . . . . . . . . . . . . . . . . . 13

ext2fs_mark_changed . . . . . . . . . . . . . . . . . . . . . . . . . 6 ext2fs_mark_ib_dirty. . . . . . . . . . . . . . . . . . . . . . . . 6

                                                ext2fs_mark_inode_bitmap . . . . . . . . . . . . . . . . . . 13
                                                ext2fs_mark_super_dirty . . . . . . . . . . . . . . . . . . . . 6

ext2fs_mark_valid . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 ext2fs_mkdir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 ext2fs_namei . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 ext2fs_namei_follow. . . . . . . . . . . . . . . . . . . . . . . . 12 ext2fs_native_flag . . . . . . . . . . . . . . . . . . . . . . . . . 17 ext2fs_new_block . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 ext2fs_new_dir_block. . . . . . . . . . . . . . . . . . . . . . . 11 ext2fs_new_inode . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 ext2fs_open . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

                                                ext2fs_open_inode_scan . . . . . . . . . . . . . . . . . . . . . 8
                                                ext2fs_parse_version_string . . . . . . . . . . . . . . . 18

ext2fs_read_bb_FILE. . . . . . . . . . . . . . . . . . . . . . . . 15 ext2fs_read_bb_inode. . . . . . . . . . . . . . . . . . . . . . . 15 ext2fs_read_bitmaps. . . . . . . . . . . . . . . . . . . . . . . . 13

                                                ext2fs_read_block_bitmap . . . . . . . . . . . . . . . . . . 13

ext2fs_read_dir_block . . . . . . . . . . . . . . . . . . . . . 11 ext2fs_read_inode . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

                                                ext2fs_read_inode_bitmap . . . . . . . . . . . . . . . . . . 13
                                                ext2fs_resize_block_bitmap . . . . . . . . . . . . . . . . 15
                                                ext2fs_resize_generic_bitmap . . . . . . . . . . . . . . 14
                                                ext2fs_resize_inode_bitmap . . . . . . . . . . . . . . . . 14

Function and Type Index 21

ext2fs_set_dir_block. . . . . . . . . . . . . . . . . . . . . . . 16

ext2fs_set_inode_callback . . . . . . . . . . . . . . . . . . 9
ext2fs_swap_group_desc . . . . . . . . . . . . . . . . . . . . 17

ext2fs_swap_inode . . . . . . . . . . . . . . . . . . . . . . . . . . 17 ext2fs_swap_super . . . . . . . . . . . . . . . . . . . . . . . . . . 17 ext2fs_test_bb_dirty. . . . . . . . . . . . . . . . . . . . . . . . 6

ext2fs_test_block_bitmap . . . . . . . . . . . . . . . . . . 13

ext2fs_test_changed . . . . . . . . . . . . . . . . . . . . . . . . . 6 ext2fs_test_ib_dirty. . . . . . . . . . . . . . . . . . . . . . . . 6

ext2fs_test_inode_bitmap . . . . . . . . . . . . . . . . . . 13

ext2fs_test_valid . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 ext2fs_unlink . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

                                                ext2fs_unmark_block_bitmap . . . . . . . . . . . . . . . . 13
                                                ext2fs_unmark_inode_bitmap . . . . . . . . . . . . . . . . 13

ext2fs_unmark_valid . . . . . . . . . . . . . . . . . . . . . . . . . 6

                                                ext2fs_update_bb_inode . . . . . . . . . . . . . . . . . . . . 15

ext2fs_write_bitmaps. . . . . . . . . . . . . . . . . . . . . . . 13

                                                ext2fs_write_block_bitmap . . . . . . . . . . . . . . . . . 13
                                                ext2fs_write_dir_block . . . . . . . . . . . . . . . . . . . . 11

ext2fs_write_inode . . . . . . . . . . . . . . . . . . . . . . . . . . 8

                                                ext2fs_write_inode_bitmap . . . . . . . . . . . . . . . . . 13

22 The EXT2FS Library (version 1.44.5)

Table of Contents

The EXT2FS Library . . . . . . . . . . . . . . . . . . . . . . . . . 1

1 Introduction to the EXT2FS Library. . . . . . . . 2

2 EXT2FS Library Functions . . . . . . . . . . . . . . . . . 3

2.1 Filesystem-level functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 2.1.1 Opening an ext2 filesystem . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 2.1.2 Closing and flushing out changes . . . . . . . . . . . . . . . . . . . . . . . . . 3 2.1.3 Initializing a filesystem . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 2.1.4 Filesystem flag functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

2.2 File I/O Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 2.2.1 File handle manipulation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 2.2.2 Reading and writing data . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 2.2.3 Changing the file offset . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 2.2.4 Getting the file size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

2.3 Inode Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 2.3.1 Reading and writing inodes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 2.3.2 Iterating over inodes in a filesystem . . . . . . . . . . . . . . . . . . . . . . 8 2.3.3 Iterating over blocks in an inode . . . . . . . . . . . . . . . . . . . . . . . . . 9 2.3.4 Convenience functions for Inodes. . . . . . . . . . . . . . . . . . . . . . . . 10

2.4 Directory functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 2.4.1 Directory block functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 2.4.2 Iterating over a directory . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 2.4.3 Creating and expanding directories. . . . . . . . . . . . . . . . . . . . . . 11 2.4.4 Creating and removing directory entries . . . . . . . . . . . . . . . . . 12 2.4.5 Looking up filenames. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 2.4.6 Translating inode numbers to filenames. . . . . . . . . . . . . . . . . . 13

2.5 Bitmap Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 2.5.1 Reading and Writing Bitmaps . . . . . . . . . . . . . . . . . . . . . . . . . . 13 2.5.2 Allocating Bitmaps . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 2.5.3 Freeing bitmaps . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 2.5.4 Bitmap Operations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 2.5.5 Comparing bitmaps . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 2.5.6 Modifying Bitmaps . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 2.5.7 Resizing Bitmaps . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 2.5.8 Clearing Bitmaps . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

2.6 EXT2 data abstractions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 2.6.1 Badblocks list management . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 2.6.2 Directory-block list management . . . . . . . . . . . . . . . . . . . . . . . . 15 2.6.3 Inode count functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

2.7 Byte-swapping functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 2.8 Other functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 ii The EXT2FS Library (version 1.44.5)

Concept Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19

Function and Type Index . . . . . . . . . . . . . . . . . . . . . 20